### v1

# Parameterized Controls and Isomorphic Representations of G-Expectations

Abstract
This paper establishes a fundamental isomorphism between the abstract framework of $G$-expectation and a concrete parameterized family of classical stochastic processes, providing a unified framework for nonlinear stochastic calculus under model uncertainty. We demonstrate that $G$-Brownian motion and related stochastic differential equations (SDEs) can be intrinsically represented as controlled processes $\{B_{t}^{\theta}\}_{\theta\in\mathcal{A}^{\Theta}}$, where $\theta$ denotes an admissible control process taking values in a compact convex parameter set $\Theta$ encoding volatility uncertainty.
The core contribution is the construction of a complete algebraic and analytic isomorphism $\Phi$ that preserves all operations including addition, multiplication, function composition, and limit transitions. This isomorphism provides a rigorous mathematical foundation for understanding $G$-martingales, with a distinction between symmetric martingales (fair in all volatility scenarios) and asymmetric martingales (favorable but fair only in the worst-case scenario).
We develop a comprehensive theory for solving $G$-SDEs through their parameterized representations, establishing existence and uniqueness results under standard Lipschitz and growth conditions. The framework reveals how bounded control sets preserve these classical conditions while capturing volatility uncertainty.
The parameterized control perspective offers both mathematical rigor and computational practicality, enabling the transfer of classical stochastic calculus results to the $G$-expectation framework while maintaining the essential sublinear structure of model uncertainty.
Key Words: G-Expectation, G-Brownian Motion, Parameterized Controls, Isomorphic Representations, Volatility Uncertainty, Dynamic Programming Principle
E-mail: zhaoguangqian@mail.ustc.edu.cn
1 Introduction
The quantification and management of model uncertainty—often referred to as Knightian uncertainty or ambiguity—represents a fundamental challenge across several disciplines, most notably in mathematical finance and decision theory. Traditional stochastic calculus, built upon the foundation of a single, precisely known probability measure, is inherently limited in its capacity to address situations where the model itself is imperfectly known. This limitation is particularly acute in financial markets, where volatility—the key parameter in option pricing and risk management—is notoriously difficult to estimate and is subject to significant time-varying uncertainty.
The groundbreaking framework of $G$-expectation, pioneered by Peng , provides a rigorous mathematical foundation for modeling such uncertainty. By replacing the classical linear expectation with a sublinear operator, the $G$-framework successfully captures the essence of volatility ambiguity. Within this framework, the $G$-Brownian motion $B_{t}$ and related stochastic differential equations ($G$-SDEs) have been defined, with their quadratic variation $\langle B\rangle_{t}$ evolving within a prescribed interval $\underline{\sigma}^{2}t,\overline{\sigma}^{2}t]$, naturally encoding the uncertainty in volatility. This leads to a rich theory of nonlinear stochastic calculus .
A profound insight emerging from this framework is the intrinsic connection between $G$-expectations and stochastic control theory. Specifically, the $G$-expectation can be naturally represented as the value function of a stochastic control problem:
$\hat{\mathbb{E}}[\varphi(B_{T})]=\sup_{\theta\in\mathcal{A}_{0,T}^{\Theta}}\mathbb{E}_{\mathbb{P}}\left[\varphi\left(\int_{0}^{T}\theta_{s}dW_{s}\right)\right],$
where the control process $\theta_{s}$ takes values in a compact set $\Theta$ representing the volatility uncertainty. This representation naturally leads to a Hamilton-Jacobi-Bellman (HJB) equation through the Dynamic Programming Principle (DPP):
$\partial_{t}v+G(D_{x}^{2}v)=0,\quad v(T,x)=\varphi(x),$
where the nonlinearity $G(A)=\frac{1}{2}\sup_{\gamma\in\Theta}\text{tr}(A\gamma\gamma^{\top})$ encodes the uncertainty set. The DPP manifests infinitesimally as:
$v(t,x)=\sup_{\theta\in\mathcal{A}_{t,t+h}^{\Theta}}\mathbb{E}_{\mathbb{P}}\left[v\left(t+h,x+\int_{t}^{t+h}\theta_{s}dW_{s}\right)\right].$
Despite this elegant connection, the abstract nature of the $G$-expectation can present a barrier to intuitive understanding and practical application. A central question thus arises:
Can we establish a more explicit and operational connection between the abstract $G$-expectation space and the concrete framework of parameterized controlled processes?
This paper provides a definitive affirmative answer by constructing a complete isomorphism between these two domains. We demonstrate that the entire $G$-expectation framework admits a rigorous correspondence to a family of classical stochastic processes controlled by volatility parameters. The core idea is to represent a $G$-Brownian motion path as a family of classical Itô processes:
$\{B_{t}^{\theta}\}_{\theta\in\mathcal{A}^{\Theta}},\quad\text{where}\quad B_{t}^{\theta}=\int_{0}^{t}\theta_{s}dW_{s}.$
The sublinear $G$-expectation is then revealed to be the supremum over all these classical scenarios, providing a stochastic control interpretation that complements the PDE perspective.
This isomorphic correspondence, denoted by $\Phi$, is shown to preserve the entire algebraic and analytic structure—including addition, multiplication, function composition, and limit transitions—thereby bridging the abstract nonlinear theory with the well-understood classical theory.
Contributions and Outline: The main contributions of this paper are as follows:
  • We construct a precise isomorphism $\Phi$ between the $G$-expectation space and a space of parameterized controlled processes, providing a concrete representation that complements the HJB equation characterization (Section 3).
  • We develop a comprehensive theory for $G$-SDEs, proving existence and uniqueness of solutions through their parameterized representations, thereby connecting the SDE solutions with the dynamic programming framework (Section 4).
  • Through this isomorphism, we characterize the fundamental structure of $G$-martingales, introducing a distinction between symmetric martingales (fair in all volatility scenarios) and asymmetric martingales (favorable but fair only in the worst-case scenario) (Sections 5 .
  • We resolve subtle definability difficulties associated with conditional $G$-expectations, clarifying the interplay between supremum operations and conditional expectations under model uncertainty.
The paper is organized as follows: Section 2 lays the theoretical foundation, formalizing $G$-expectation as a value function and deriving its connection to dynamic programming and nonlinear PDEs. Section 3 introduces the fundamental isomorphism, the central concept of this work. Sections 4 and 5 delve into the solutions of $G$-SDEs and the properties of $G$-martingales from this new perspective, respectively. Finally, Section 6 concludes.
Our work not only deepens the theoretical understanding of the $G$-framework but also unlocks new computational pathways. By translating complex nonlinear problems into families of more tractable classical ones, this parameterized control perspective offers a powerful tool for both theoretical analysis and practical computation in the presence of model uncertainty, while maintaining a natural connection to the HJB and dynamic programming framework.
2 Theoretical Foundation
2.1 G-Expectation as a Value Function
The fundamental insight is that G-expectation can be represented as the value function of a stochastic control problem. For a G-Brownian motion $B_{t}$ and terminal payoff $\varphi\in C_{b}(\mathbb{R}^{d})$, we have:
$\hat{\mathbb{E}}[\varphi(B_{T})]=\sup_{\theta\in\mathcal{A}^{\Theta}_{0,T}}\mathbb{E}_{\mathbb{P}}\left[\varphi\left(\int_{0}^{T}\theta_{s}dW_{s}\right)\right]$ (2.1)
where:
  • $\mathcal{A}^{\Theta}_{0,T}$ denotes admissible controls taking values in $\Theta$
  • $W_{t}$ is a classical Brownian motion under reference measure $\mathbb{P}$
  • $\Theta=[\underline{\sigma}^{2}I_{d},\overline{\sigma}^{2}I_{d}]\subset S^{d}_{+}$ models volatility uncertainty
This equivalence reveals that:
  1. 1.
    Volatility ambiguity is characterized by the control set $\Theta$
  2. 2.
    G-Brownian paths emerge as controlled processes: $B_{t}^{\theta}=\int_{0}^{t}\theta_{s}dW_{s}$
2.2 Dynamic Programming and Nonlinear PDEs
The control-theoretic perspective leads to a Hamilton-Jacobi-Bellman (HJB) equation. Define the value function:
$v(t,x)=\sup_{\theta\in\mathcal{A}^{\Theta}_{t,T}}\mathbb{E}_{\mathbb{P}}\left[\varphi\left(x+\int_{t}^{T}\theta_{s}dW_{s}\right)\right]$ (2.2)
It satisfies the nonlinear parabolic PDE:
$\begin{cases}\partial_{t}v+G(D_{x}^{2}v)=0,&(t,x)\in[0,T)\times\mathbb{R}^{d}\\ v(T,x)=\varphi(x)\end{cases}$ (2.3)
where the nonlinearity $G$ encodes the uncertainty set:
$G(A)=\frac{1}{2}\sup_{\gamma\in\Theta}\text{tr}(A\gamma\gamma^{\top}),\quad A\in S^{d}$ (2.4)
This PDE is the infinitesimal manifestation of the Dynamic Programming Principle (DPP):
$v(t,x)=\sup_{\theta\in\mathcal{A}^{\Theta}_{t,t+h}}\mathbb{E}_{\mathbb{P}}\left[v\left(t+h,x+\int_{t}^{t+h}\theta_{s}dW_{s}\right)\right]$ (2.5)
3 The Fundamental Isomorphism: G-Brownian motion as Parameterized Controlled Processes
The central insight of this work is establishing a natural isomorphism between elements in the G-expectation space and parameterized families of controlled processes. Specifically:
  1. 1.
    G-Brownian motion as controlled paths: For each volatility control process $\theta\in\Theta$, we define:
    $B_{t}^{\theta}=\int_{0}^{t}\theta_{s}dW_{s}$
    with quadratic variation correctly given by:
    $d\langle B^{\theta}\rangle_{t}^{ij}=(\theta_{t}\theta_{t}^{\top})^{ij}dt$
  2. 2.
    G-SDE solutions as parameterized families: For a general G-SDE:
    $dX_{t}=b(X_{t})dt+h_{ij}(X_{t})d\langle B\rangle_{t}^{ij}+\sigma_{j}(X_{t})dB_{t}^{j}$
    the solution is intrinsically represented by the parameterized family:
    $X_{t}^{\theta}=X_{0}+\int_{0}^{t}b(X_{s}^{\theta})ds+\int_{0}^{t}h_{ij}(X_{s}^{\theta})(\theta_{s}\theta_{s}^{\top})^{ij}ds+\int_{0}^{t}\sigma_{j}(X_{s}^{\theta})\theta_{s}^{j}dW_{s}^{j}$
  3. 3.
    The isomorphism: There exists a bijective correspondence:
    $\begin{array}[]{c}\text{G-expectation space}\\ \Updownarrow\\ \left\{X^{\theta}:\theta\in\Theta\right\}\end{array}$
    where the parameterized family $\{X^{\theta}\}_{\theta\in\Theta}$ is the G-SDE solution
This isomorphism reveals that:
  • G-processes are intrinsically controlled trajectories parameterized by volatility uncertainty
  • The G-expectation framework provides a geometric encoding of model uncertainty
  • Solutions to G-SDEs emerge as universal families of classical SDE solutions
3.1 Classical G-Expectation and G-Brownian Motion
3.1.1 G-Normal Distribution
The cornerstone of the theory is the G-normal distribution . Let $\xi$ be a random variable. It is said to follow a G-normal distribution, denoted $\xi\sim\mathcal{N}(0,[\underline{\sigma}^{2},\overline{\sigma}^{2}])$, if for any $\phi\in C_{b}(\mathbb{R})$, its sublinear expectation is defined by the following nonlinear heat equation:
$\mathbb{E}[\phi(\xi)]=u(1,0),$ (3.1)
where $u(t,x)$ is the unique viscosity solution to the following nonlinear partial differential equation (PDE):
$\frac{\partial u}{\partial t}$ $=G\left(\frac{\partial^{2}u}{\partial x^{2}}\right),\quad(t,x)\in(0,\infty)\times\mathbb{R},$ (3.2)
$u(0,x)$ $=\phi(x),$ (3.3)
with the nonlinear generator $G:\mathbb{R}\to\mathbb{R}$ given by:
$G(\alpha)=\frac{1}{2}\sup_{\sigma\in[\underline{\sigma},\overline{\sigma}]}(\sigma^{2}\alpha)=\frac{1}{2}\left(\overline{\sigma}^{2}\alpha^{+}-\underline{\sigma}^{2}\alpha^{-}\right).$ (3.4)
Here, $\alpha^{+}=\max(\alpha,0)$ and $\alpha^{-}=-\min(\alpha,0)$. This distribution is normal in the sense that it is centered and stable under scaling and addition of independent random variables, yet it is not linear due to the sublinearity of $G$.
A particularly revealing special case arises when the uncertainty set collapses to a single point, i.e., $\underline{\sigma}=\overline{\sigma}=\sigma_{0}>0$. In this case, the generator becomes linear, $G(\alpha)=\frac{1}{2}\sigma_{0}^{2}\alpha$, and the G-normal distribution reduces to the classical normal distribution $\mathcal{N}(0,\sigma_{0}^{2})$.
3.1.2 G-Brownian Motion
[…]
  1. (i)
    $B_{0}=0$.
  2. (ii)
    For any $t,s\geq 0$, the increment $B_{t+s}-B_{t}$ is identically distributed to $B_{s}$, and $B_{t+s}-B_{t}$ is independent of $(B_{t_{1}},B_{t_{2}},\dots,B_{t_{n}})$ for any $n\in\mathbb{N}$ and $0\leq t_{1}\leq\dots\leq t_{n}\leq t$.
  3. (iii)
    $\lim_{t\downarrow 0}\hat{E}[|B_{t}|^{3}]t^{-1}=0$.
Moreover, if $\hat{E}[B_{t}]=\hat{E}[-B_{t}]=0$ for all $t$, then $(B_{t})_{t\geq 0}$ is called a symmetric G-Brownian motion.
In the context of G-expectation, the increments of G-Brownian motion follow a G-normal distribution. Specifically, for any $t,s\geq 0$, $B_{t+s}-B_{t}\sim\mathcal{N}(0,[\underline{\sigma}^{2}s,\overline{\sigma}^{2}s])$, where $\mathcal{N}(0,[\underline{\sigma}^{2},\overline{\sigma}^{2}])$ denotes the G-normal distribution with uncertainty interval $\underline{\sigma}^{2},\overline{\sigma}^{2}]$.
The quadratic variation process $\langle B\rangle_{t}$ of G-Brownian motion is a continuous, non-decreasing process such that ,
$\underline{\sigma}^{2}t\leq\langle B\rangle_{t}\leq\overline{\sigma}^{2}t.$
This process captures the accumulated uncertainty over time.
The corresponding infinitesimal generator for a function $\phi$ is given by the same nonlinear operator $G$:
$\lim_{t\downarrow 0}\frac{\mathbb{E}[\phi(x+B_{t})]-\phi(x)}{t}=G(\phi^{\prime\prime}(x)).$ (3.5)
3.2 A Generalized Construction via Random Variables
This generator governs the evolution of expectations under the G-Brownian motion. The drift term is zero, reflecting the martingale property of $B_{t}$, and the diffusion term is given by the supremum over constant parameters. This establishes the foundational link: classical G-Brownian motion corresponds to a constant diffusion coefficient uncertainty set $\Theta=[\underline{\sigma},\overline{\sigma}]$.
Definition 3.1 (Construction via Supremum over Linear Expectations).
Let $(\Omega,\mathcal{F},P)$ be a probability space sufficient to support a standard Brownian motion $\{W_{t}\}_{t\geq 0}$. For a given compact, convex parameter set $\Lambda\subset\mathbb{R}^{n}$, we define two fundamental random objects:
1. Max-Distribution: For a parameter $\mu\in\Lambda$, define the random variable $\eta(\mu)=\mu$. The associated nonlinear expectation $\mathbb{E}[\eta]$ is defined for a test function $\phi$ as:
$\mathbb{E}[\phi(\eta)]:=\sup_{\mu\in\Lambda}E_{P}[\phi(\eta(\mu))].$ (3.6)
This object captures the pure parameter uncertainty in its most extreme form.
2. G-Normal Distribution: For a parameter $\sigma\in\Lambda$, define the random variable $X(\sigma)=\sigma W_{1}$ (which is classically normal distributed, $X(\sigma)\sim N(0,\sigma^{2})$). The associated nonlinear expectation $\mathbb{E}[X]$ is defined as:
$\mathbb{E}[\phi(X)]:=\sup_{\sigma\in\Lambda}E_{P}[\phi(X(\sigma))]=\sup_{\sigma\in\Lambda}E_{P}[\phi(\sigma W_{1})].$ (3.7)
This constitutes a G-normal distribution with the uncertainty set $\Lambda$.
3. Quasi G-Brownian Motion: We define a process $\{B_{t}\}_{t\geq 0}$ by its action on test functions. For any $\phi\in C_{l,Lip}(\mathbb{R})$, its nonlinear expectation is given by:
$\mathbb{E}[\phi(B_{t})]:=\sup_{\sigma\in\Lambda}E_{P}[\phi(\sigma W_{t})].$ (3.8)
Furthermore, for a process of the form $H_{t}=bB_{t}+mt$ (where the drift $m$ is also subject to uncertainty within a convex set), the expectation is defined as:
$\mathbb{E}[\phi(H_{t})]:=\sup_{(b,m)\in\Lambda^{\prime}}E_{P}[\phi(bW_{t}+mt)],$ (3.9)
where $\Lambda^{\prime}$ is a compact convex set in $\mathbb{R}^{2}$.
These constructions are defined on a filtered probability space represented by the quadruple:
$(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\mathbb{E})$ (3.10)
where:
  • $\Omega$ is the sample space, representing the set of all possible outcomes;
  • $\mathcal{F}$ is a $\sigma$-algebra on $\Omega$, representing the set of all measurable events;
  • $(\mathcal{F}_{t})_{t\geq 0}$ is a filtration, i.e., an increasing family of sub-$\sigma$-algebras of $\mathcal{F}$ satisfying $\mathcal{F}_{s}\subseteq\mathcal{F}_{t}$ for all $0\leq s\leq t$, which represents the accumulation of information over time;
  • $\mathbb{E}$ is the sublinear expectation operator defined in equations (3.6) and (3.7), which generalizes the classical linear expectation to account for model uncertainty.
Remark 3.1.
The construction of the ”Quasi G-Brownian Motion” and the related objects herein represents the core methodological approach of this paper. It is intended to schematically illustrate the central idea of handling model uncertainty via the supremum over linear expectations. It is important to note that the process $\{B_{t}\}$ defined here is for methodological demonstration purposes; while it shares the same one-dimensional distributions with the strict mathematical G-Brownian motion, their joint distributions are entirely different, and thus they are not mathematically equivalent.
Definition 3.2 (General Construction of Sublinear Expectation via Model Uncertainty).
Let $(\Omega,\mathcal{F},P)$ be a classical probability space. Let $\Theta$ be a non-empty, compact, and convex parameter set representing the domain of model uncertainty.
Consider a generalparameter-controlled random variable $X(\theta)$, where $\theta\in\Theta$ is the parameter that dictates the model specification. The associated sublinear expectation $\mathbb{E}[\cdot]$ is defined for a test function $\phi$ (typically in the space of Lipschitz continuous functions $C_{l,Lip}(\mathbb{R}^{d})$) by the following operation:
$\mathbb{E}[\phi(X)]:=\sup_{\theta\in\Theta}E_{P}[\phi(X(\theta))].$ (3.11)
This means the operator $\mathbb{E}$ acts on the random variable $X$ by a two-step procedure:
1. For a fixed parameter $\theta\in\Theta$, compute the classical linear expectation $E_{P}$ of the random variable $\phi(X(\theta))$ under the probability measure $P$.
2. Subsequently, take the supremum of these resulting real numbers over all admissible parameters $\theta$ in the uncertainty set $\Theta$.
This operator is defined on the sublinear expectation space:
$(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\mathbb{E})$ (3.12)
where:
  • $\Omega$ is the sample space;
  • $\mathcal{F}$ is a $\sigma$-algebra on $\Omega$;
  • $(\mathcal{F}_{t})_{t\geq 0}$ is a filtration representing the information flow over time;
  • $\mathbb{E}$ is the sublinear expectation operator defined above.
Remarks:
  • This general definition subsumes the specific constructions (e.g., Max-Distribution, G-Normal Distribution, Generalized G-Brownian Motion) presented in Definition 3.1 as special cases by choosing the parameter-controlled variable $X(\theta)$ and the uncertainty set $\Theta$ appropriately.
  • The operator $\mathbb{E}$ defined in this way is sublinear: $\mathbb{E}[X+Y]\leq\mathbb{E}[X]+\mathbb{E}[Y]$ and $\mathbb{E}[\lambda X]=\lambda\mathbb{E}[X]$ for $\lambda\geq 0$. It also preserves constants: $\mathbb{E}[c]=c$ for any constant $c$.
While the parameterized construction provides valuable insight, it is crucial to recognize its fundamental limitation: constant control sets cannot capture the path-dependent nature of true G-Brownian motion.
Consider the parameterized ”quasi G-Brownian motion” defined by:
$B_{t}^{\sigma}=\sigma W_{t},\quad\sigma\in\Lambda$ (3.13)
with nonlinear expectation:
$\mathbb{E}[\phi(B_{t})]:=\sup_{\sigma\in\Lambda}E_{P}[\phi(\sigma W_{t})]$ (3.14)
Key observations:
  1. 1.
    One-dimensional distributions match: For fixed $t$,
    $\text{Law}_{G}(B_{t})\cong\sup_{\sigma\in\Lambda}\text{Law}_{P}(\sigma W_{t})$
    meaning marginal distributions coincide with true G-Brownian motion.
  2. 2.
    Path structure diverges: The joint distributions differ fundamentally:
    $\mathbb{E}[\phi(B_{t},B_{s})]=\sup_{\sigma\in\Lambda}E_{P}[\phi(\sigma W_{t},\sigma W_{s})]\neq\hat{\mathbb{E}}[\phi(B_{t},B_{s})]$
    This occurs because constant $\sigma$ cannot capture the time-consistent volatility uncertainty of true G-Brownian motion.
  3. 3.
    The Difficulty in Defining Conditional Expectation
3.3 True G-Brownian
To overcome these limitations, we must introduce time-varying adaptive controls:
Definition 3.3 (True G-Brownian Motion via Adaptive Controls).
Let $\Theta\subset\mathbb{R}^{d\times d}$ be compact convex. Define:
$B_{t}^{\theta}=\int_{0}^{t}\theta_{s}dW_{s},\quad\theta\in\mathcal{A}^{\Theta}$ (3.15)
with quadratic variation:
$d\langle B^{\theta}\rangle_{t}=(\theta_{t}\theta_{t}^{\top})dt$ (3.16)
The genuine G-Brownian motion is then characterized by:
${\mathbb{E}}[\phi(B_{t})]=\sup_{\theta\in\mathcal{A}^{\Theta}}E_{P}[\phi(B_{t}^{\theta})]$ (3.17)
The space of admissible control processes is defined as:
$\mathcal{A}^{\Theta}=\left\{\theta:[0,\infty)\times\Omega\to\Theta\ \middle|\ \theta\text{ is }\mathcal{F}_{t}\text{-adapted and }\mathbb{E}\left[\int_{0}^{T}|\theta_{t}|^{2}dt\right]<\infty\ \forall T>0\right\}$
where $\Theta\subset\mathbb{R}^{d\times d}$ is a compact convex set representing the volatility uncertainty.
For the G-Brownian motion $B_{t}$, the conditional expectation is:
${\mathbb{E}}_{t}[\varphi(B_{T})]=\sup_{\theta\in\mathcal{A}_{t,T}^{\Theta}}\mathbb{E}_{\mathbb{P}}\left[\varphi\left(\int_{t}^{T}\theta_{s}dW_{s}\right)\middle|\mathcal{F}_{t}\right]$
This definition provides the mathematical foundation for conditional expectations in the G-framework, capturing the essence of model uncertainty through the supremum over volatility controls.
Definition 3.4 (Conditional G-Expectation Operator).
For a G-Brownian motion $\{B_{t}\}_{t\geq 0}$ and a test function $\varphi\in C_{b.Lip}(\mathbb{R}^{d})$, the conditional G-expectation operator $\hat{\mathbb{E}}_{s}[\cdot]$ is defined as:
${\mathbb{E}}_{s}[\varphi(B_{t})]=\sup_{\theta\in\mathcal{A}_{s,t}^{\Theta}}\mathbb{E}_{\mathbb{P}}\left[\varphi\left(B_{s}+\int_{s}^{t}\theta_{u}dW_{u}\right)\middle|\mathcal{F}_{s}\right]$
where:
  • $\mathcal{A}_{s,t}^{\Theta}$ denotes admissible control processes on $[s,t]$
  • $\theta_{u}\in\Theta=[\underline{\sigma},\overline{\sigma}]$ represents volatility controls
  • $W_{u}$ is a standard Brownian motion under $\mathbb{P}$
Definition 3.5 (Conditional G-Expectation for Cylinder Functions).
For any cylinder function of the form:
$\xi=\varphi(B_{t_{1}},B_{t_{2}}-B_{t_{1}},\cdots,B_{t_{n}}-B_{t_{n-1}}),\quad\varphi\in C_{b.Lip}(\mathbb{R}^{d\times n})$
with $0\leq t_{1}<t_{2}<\cdots<t_{n}\leq T$, the conditional G-expectation is defined as:
${\mathbb{E}}_{s}[\xi]=\sup_{\theta\in\mathcal{A}_{s,T}^{\Theta}}\mathbb{E}_{\mathbb{P}}\left[\varphi\left(B_{t_{1}}^{\theta},B_{t_{2}}^{\theta}-B_{t_{1}}^{\theta},\cdots,B_{t_{n}}^{\theta}-B_{t_{n-1}}^{\theta}\right)\middle|\mathcal{F}_{s}\right]$
where $B_{t}^{\theta}=\int_{0}^{t}\theta_{u}dW_{u}$ is the controlled process.
This definition coincides with the classical G-expectation definition when $s=0$:
${\mathbb{E}}[\xi]=\sup_{\theta\in\mathcal{A}_{0,T}^{\Theta}}\mathbb{E}_{\mathbb{P}}\left[\varphi\left(B_{t_{1}}^{\theta},B_{t_{2}}^{\theta}-B_{t_{1}}^{\theta},\cdots,B_{t_{n}}^{\theta}-B_{t_{n-1}}^{\theta}\right)\right]$
Theorem 3.6 (Conditional Expectation Simplification).
For any cylinder function $\xi=\varphi(B_{t_{1}},B_{t_{2}}-B_{t_{1}},\cdots,B_{t_{n}}-B_{t_{n-1}})$ and $s\leq T$, the conditional G-expectation simplifies to:
${\mathbb{E}}_{s}[\xi]=\sup_{\theta\in\mathcal{A}_{s,T}^{\Theta}}\mathbb{E}_{\mathbb{P}}\left[\varphi\left(x_{1},x_{2},\cdots,x_{k-1},B_{t_{k}}^{\theta}-B_{t_{k-1}}^{\theta},\cdots,B_{t_{n}}^{\theta}-B_{t_{n-1}}^{\theta}\right)\right]\bigg{|}_{x_{i}=B_{t_{i}}}$
where for $t_{k-1}\leq s$$t_{k}>s$, we treat $B_{t_{i}}$ as constants $x_{i}$ in the expectation.
Note that our definition of conditional expectation here is completely consistent with the classical definition of G-conditional expectation under the definition of cylindrical Brownian motion, so we can similarly extend it to its completion space.
Theorem 3.7 (Fundamental Properties).
The conditional G-expectation operator satisfies:
  1. 1.
    Sublinearity: ${\mathbb{E}}_{s}[X+Y]\leq{\mathbb{E}}_{s}[X]+{\mathbb{E}}_{s}[Y]$
  2. 2.
    Positive Homogeneity: ${\mathbb{E}}_{s}[\lambda X]=\lambda{\mathbb{E}}_{s}[X]$ for $\lambda\geq 0$
  3. 3.
    Preservation of Constants: ${\mathbb{E}}_{s}[c]=c$ for any constant $c$
  4. 4.
    Monotonicity: If $X\leq Y$ , then ${\mathbb{E}}_{s}[X]\leq{\mathbb{E}}_{s}[Y]$
Theorem 3.8 (Tower Property).
For $s\leq t\leq u$, the conditional expectation satisfies:
${\mathbb{E}}_{s}\left[{\mathbb{E}}_{t}[X]\right]={\mathbb{E}}_{s}[X]$
This ensures the time consistency of the G-expectation framework.
3.4 The Complete Isomorphism
To establish a full isomorphism, we must introduce time-varying adaptive controls.
3.4.1 Adaptive Control Representation
Definition 3.9 (True G-Brownian Motion via Adaptive Controls).
Let $\Theta\subset\mathbb{R}$ be compact convex. Define:
$B_{t}^{\theta}=\int_{0}^{t}\theta_{s}dW_{s},\quad\theta\in\mathcal{A}^{\Theta}$ (3.18)
with quadratic variation:
$d\langle B^{\theta}\rangle_{t}=\theta_{t}^{2}dt$ (3.19)
The genuine G-Brownian motion is characterized by:
$\hat{\mathbb{E}}[\phi(B_{t})]=\sup_{\theta\in\mathcal{A}^{\Theta}}\mathbb{E}_{\mathbb{P}}[\phi(B_{t}^{\theta})]$ (3.20)
3.4.2 Algebraic Structure Preservation
The isomorphism preserves all algebraic operations:
Theorem 3.10 (Algebraic Isomorphism).
The mapping $\Phi:X\mapsto\{X^{\theta}\}_{\theta\in\mathcal{A}^{\Theta}}$ preserves:
  1. 1.
    Addition: $\Phi(X+Y)=\Phi(X)+\Phi(Y)$
  2. 2.
    Scalar Multiplication: $\Phi(\lambda X)=\lambda\Phi(X)$
  3. 3.
    Multiplication: $\Phi(X\cdot Y)=\Phi(X)\cdot\Phi(Y)$
  4. 4.
    Function Composition: $\Phi(\phi(X))=\phi(\Phi(X))$
The preservation follows directly from the pathwise definition:
$(X+Y)^{\theta}$ $=X^{\theta}+Y^{\theta}$
$(\lambda X)^{\theta}$ $=\lambda X^{\theta}$
$(X\cdot Y)^{\theta}$ $=X^{\theta}\cdot Y^{\theta}$
$\phi(X)^{\theta}$ $=\phi(X^{\theta})$
Fourth, we define the multivariate composition in a similar manner. Building upon the correspondence established via G-Brownian motion, we derive a corresponding relationship in the space spanned by these two elements, while ensuring that the mapping constitutes a bijection.
3.4.3 Analytic Structure Preservation
The isomorphism also preserves analytic properties:
Theorem 3.11 (Analytic Isomorphism).
The mapping $\Phi$ preserves:
  1. 1.
    Expectation: $\hat{\mathbb{E}}[X]=\sup_{\theta}\mathbb{E}_{\mathbb{P}}[X^{\theta}]$
  2. 2.
    Conditional Expectation: $\hat{\mathbb{E}}_{t}[X]=\sup_{\theta}\mathbb{E}_{\mathbb{P}}[X^{\theta}|\mathcal{F}_{t}]$
  3. 3.
    Convergence: $X_{n}\to X$ in $L_{G}^{p}$ iff $X_{n}^{\theta}\to X^{\theta}$ in $L^{p}$ uniformly over $\theta$
3.5 Stochastic Calculus Equivalence
The isomorphism extends to differential operations:
3.5.1 Itô Calculus Preservation
Theorem 3.12 (Itô Calculus Isomorphism).
For G-Itô processes:
$dX_{t}$ $=b_{t}dt+h_{t}d\langle B\rangle_{t}+\sigma_{t}dB_{t}$
$dX_{t}^{\theta}$ $=b_{t}^{\theta}dt+h_{t}^{\theta}\theta_{t}^{2}dt+\sigma_{t}^{\theta}\theta_{t}dW_{t}$
we have:
  1. 1.
    $\Phi(dB_{t})=\{\theta_{t}dW_{t}\}$
  2. 2.
    $\Phi(\langle B\rangle_{t})=\left\{\int_{0}^{t}\theta_{s}^{2}ds\right\}$
  3. 3.
    $\Phi\left(\int_{0}^{t}H_{s}dB_{s}\right)=\left\{\int_{0}^{t}H_{s}^{\theta}\theta_{s}dW_{s}\right\}$
3.5.2 Itô Calculus Isomorphism (Integral Form)
Proof.
We prove each component using integral representations and L²-convergence:
Part 1: $\Phi\left(\int_{0}^{t}dB_{s}\right)=\left\{\int_{0}^{t}\theta_{s}dW_{s}\right\}$
By definition of the isomorphism:
$\Phi\left(\int_{0}^{t}dB_{s}\right)=\Phi(B_{t})=\left\{B_{t}^{\theta}\right\}_{\theta\in\mathcal{A}^{\Theta}}=\left\{\int_{0}^{t}\theta_{s}dW_{s}\right\}_{\theta\in\mathcal{A}^{\Theta}}$
This holds identically as:
$B_{t}=\int_{0}^{t}dB_{s},\quad B_{t}^{\theta}=\int_{0}^{t}\theta_{s}dW_{s}$
Part 2: $\Phi\left(\langle B\rangle_{t}\right)=\left\{\int_{0}^{t}\theta_{s}^{2}ds\right\}$
The quadratic variation is defined as the L²-limit:
$\langle B\rangle_{t}=\lim_{\|\Pi\|\to 0}\sum_{k=1}^{n}(B_{t_{k}}-B_{t_{k-1}})^{2}\quad\text{in }L_{G}^{2}$
Applying $\Phi$:
$\Phi(\langle B\rangle_{t})=\Phi\left(\lim_{\|\Pi\|\to 0}\sum_{k=1}^{n}(B_{t_{k}}-B_{t_{k-1}})^{2}\right)$
Since $\Phi$ preserves L²-convergence:
$\Phi(\langle B\rangle_{t})=\left\{\lim_{\|\Pi\|\to 0}\sum_{k=1}^{n}(B_{t_{k}}^{\theta}-B_{t_{k-1}}^{\theta})^{2}\right\}_{\theta\in\mathcal{A}^{\Theta}}\quad\text{in }L^{2}$
For each fixed $\theta$, by Itô isometry:
$\lim_{\|\Pi\|\to 0}\sum_{k=1}^{n}\left(\int_{t_{k-1}}^{t_{k}}\theta_{s}dW_{s}\right)^{2}=\int_{0}^{t}\theta_{s}^{2}ds\quad\text{in }L^{2}(\Omega,\mathbb{P})$
Thus:
$\Phi(\langle B\rangle_{t})=\left\{\int_{0}^{t}\theta_{s}^{2}ds\right\}_{\theta\in\mathcal{A}^{\Theta}}$
Part 3: $\Phi\left(\int_{0}^{t}H_{s}dB_{s}\right)=\left\{\int_{0}^{t}H_{s}^{\theta}\theta_{s}dW_{s}\right\}$
The G-Itô integral is defined as:
$\int_{0}^{t}H_{s}dB_{s}=\lim_{\|\Pi\|\to 0}\sum_{k=1}^{n}H_{t_{k-1}}(B_{t_{k}}-B_{t_{k-1}})\quad\text{in }L_{G}^{2}$
Applying $\Phi$:
$\Phi\left(\int_{0}^{t}H_{s}dB_{s}\right)=\left\{\lim_{\|\Pi\|\to 0}\sum_{k=1}^{n}H_{t_{k-1}}^{\theta}(B_{t_{k}}^{\theta}-B_{t_{k-1}}^{\theta})\right\}_{\theta\in\mathcal{A}^{\Theta}}$
For each fixed $\theta$, the classical Itô integral satisfies:
$\lim_{\|\Pi\|\to 0}\sum_{k=1}^{n}H_{t_{k-1}}^{\theta}\int_{t_{k-1}}^{t_{k}}\theta_{s}dW_{s}=\int_{0}^{t}H_{s}^{\theta}\theta_{s}dW_{s}\quad\text{in }L^{2}(\Omega,\mathbb{P})$
Thus:
$\Phi\left(\int_{0}^{t}H_{s}dB_{s}\right)=\left\{\int_{0}^{t}H_{s}^{\theta}\theta_{s}dW_{s}\right\}_{\theta\in\mathcal{A}^{\Theta}}$
Consistency with G-Itô Formula
Consider the integral form of G-Itô formula for $f(B_{t})$:
$f(B_{t})-f(0)=\int_{0}^{t}\partial_{x}f(B_{s})dB_{s}+\frac{1}{2}\int_{0}^{t}\partial_{xx}^{2}f(B_{s})d\langle B\rangle_{s}$
Applying $\Phi$:
$\Phi(f(B_{t})-f(0))=\left\{f(B_{t}^{\theta})-f(0)\right\}_{\theta\in\mathcal{A}^{\Theta}}$
By classical Itô formula for each $\theta$:
$f(B_{t}^{\theta})-f(0)=\int_{0}^{t}\partial_{x}f(B_{s}^{\theta})\theta_{s}dW_{s}+\frac{1}{2}\int_{0}^{t}\partial_{xx}^{2}f(B_{s}^{\theta})\theta_{s}^{2}ds$
Thus:
$\Phi\left(\int_{0}^{t}\partial_{x}f(B_{s})dB_{s}+\frac{1}{2}\int_{0}^{t}\partial_{xx}^{2}f(B_{s})d\langle B\rangle_{s}\right)=\left\{\int_{0}^{t}\partial_{x}f(B_{s}^{\theta})\theta_{s}dW_{s}+\frac{1}{2}\int_{0}^{t}\partial_{xx}^{2}f(B_{s}^{\theta})\theta_{s}^{2}ds\right\}_{\theta\in\mathcal{A}^{\Theta}}$
This confirms the isomorphism preserves the integral structure of Itô calculus in L²-sense. ∎
3.6 Visualizing the Isomorphism: From G-SDE to Parameterized Processes
[…]
The isomorphism $\Phi$ establishes a complete correspondence between the G-expectation space and the parameterized classical space:
with the mapping defined as:
$\Phi(X)=\{X^{\theta}\}_{\theta\in\Theta}$
3.7 Algebraic and Analytic Structure Preservation
The isomorphism preserves all algebraic and analytic operations:
Operation G-Expectation Space Parameterized Space
Addition $X+Y$ $\Phi(X)+\Phi(Y)=\{X^{\theta}+Y^{\theta}\}$
Multiplication $X\cdot Y$ $\Phi(X)\cdot\Phi(Y)=\{X^{\theta}\cdot Y^{\theta}\}$
Scalar Multiplication $\lambda X$ $\lambda\Phi(X)=\{\lambda X^{\theta}\}$
Function Composition $\phi(X)$ $\phi(\Phi(X))=\{\phi(X^{\theta})\}$
Limit Operation $\lim\limits_{n\to\infty}X_{n}$ $\lim\limits_{n\to\infty}\Phi(X_{n})=\{\lim\limits_{n\to\infty}X_{n}^{\theta}\}$
Stochastic Integral $\int_{0}^{t}H_{s}dB_{s}$ $\left\{\int_{0}^{t}H_{s}^{\theta}\theta_{s}dW_{s}\right\}$
Quadratic Variation $\langle B\rangle_{t}$ $\left\{\int_{0}^{t}\theta_{s}^{2}ds\right\}$
The sublinear expectation operators are related through the isomorphism:
G-Expectation Space Parameterized Space
$\hat{\mathbb{E}}[X]$ $\sup\limits_{\theta\in\Theta}\mathbb{E}_{\mathbb{P}}[X^{\theta}]$
$\hat{\mathbb{E}}_{t}[X]$ $\sup\limits_{\theta\in\Theta}\mathbb{E}_{\mathbb{P}}[X^{\theta}\mid\mathcal{F}_{t}]$
  1. 1.
    Intrinsic Nature of Expectations:
    • $\hat{\mathbb{E}}$ is the intrinsic sublinear expectation operator of the G-expectation space
    • $\mathbb{E}_{\mathbb{P}}$ is the intrinsic linear expectation operator of the classical probability space
    • The supremum operation $\sup_{\theta}$ is the bridge connecting these intrinsic structures
  2. 2.
    sublinear expectation operators:
    $\hat{\mathbb{E}}[X]=\sup_{\theta\in\Theta}\mathbb{E}_{\mathbb{P}}[X^{\theta}]$
    This formula does not define $\hat{\mathbb{E}}$, but rather expresses how the intrinsic G-expectation can be computed using the parameterized classical expectations.
  3. 3.
    Analytic Continuity: The isomorphism preserves convergence properties:
    $X_{n}\to X\text{ in }{L^{p}_{\mathbb{G}}}\iff X_{n}^{\theta}\to X^{\theta}\text{ in }L^{p}\text{ uniformly over }\theta$
This isomorphism provides:
  • A computational bridge from nonlinear to linear stochastic calculus
  • A theoretical foundation for model uncertainty quantification
  • An operational framework for solving G-SDEs through parameterized classical SDEs
4 How to Solve G-Stochastic Differential Equations
Consider a classical SDE parameterized by $\theta\in\Theta$:
$dX_{t}^{\theta}=b(t,X_{t}^{\theta})dt+\sigma(t,X_{t}^{\theta})\theta_{t}dW_{t}$ (4.1)
where:
  • $W_{t}$ is a standard Brownian motion
  • $\theta_{t}\in\Theta$ is an adapted control process
  • $\Theta\subset\mathbb{R}$ is compact and convex (bounded controls)
Under standard Lipschitz and growth conditions on $b$ and $\sigma$, for each $\theta\in\mathcal{A}^{\Theta}$, there exists a unique strong solution $X_{t}^{\theta}$.
We reformulate the classical SDE as a G-SDE driven by G-Brownian motion $B_{t}$:
The classical SDE (4.1) is equivalent to the following G-SDE:
$dX_{t}=b(t,X_{t})dt+\sigma(t,X_{t})dB_{t}$ (4.2)
where $B_{t}$ is a G-Brownian motion
Theorem 4.1 (Existence and Uniqueness for G-SDE).
Under bounded controls $\theta_{t}\in\Theta$ and standard Lipschitz conditions:
$|b(t,x)-b(t,y)|+|\sigma(t,x)-\sigma(t,y)|$ $\leq K|x-y|$
$|b(t,x)|^{2}+|\sigma(t,x)|^{2}$ $\leq K(1+|x|^{2})$
the G-SDE (4.2) admits a unique solution in the quasi-sure sense:
$X_{t}=\{X_{t}^{\theta}\}_{\theta\in\mathcal{A}^{\Theta}}$
Proof.
The proof follows from:
  1. 1.
    For each fixed $\theta\in\mathcal{A}^{\Theta}$, the classical SDE has a unique strong solution $X_{t}^{\theta}$
  2. 2.
    The boundedness of $\Theta$ ensures uniform estimates:
    $\sup_{\theta\in\mathcal{A}^{\Theta}}\mathbb{E}\left[\sup_{0\leq s\leq t}|X_{s}^{\theta}|^{2}\right]<\infty$
  3. 3.
    The solution family $\{X_{t}^{\theta}\}$ satisfies the consistency conditions for G-expectation
where $\theta_{t}\in\Theta$ is an adapted control process with $\Theta\subset\mathbb{R}$ compact and convex.
4.1 Existence and Uniqueness Theorem
Theorem 4.2 (Classical SDE Solution).
Assume the coefficients satisfy:
  1. 1.
    Lipschitz continuity: For all $x,y\in\mathbb{R}$, $t\in[0,T]$,
    $|b(t,x)-b(t,y)|+|\sigma(t,x)-\sigma(t,y)|\leq K|x-y|$ (4.3)
  2. 2.
    Linear growth: For all $x\in\mathbb{R}$, $t\in[0,T]$,
    $|b(t,x)|+|\sigma(t,x)|\leq K(1+|x|)$ (4.4)
  3. 3.
    Bounded controls: $\theta_{t}\in\Theta$ with $\Theta$ compact, so
    $\exists M>0:|\theta_{t}|\leq M\quad\forall t,\omega$ (4.5)
Then for each $\theta\in\mathcal{A}^{\Theta}$, there exists a unique strong solution $X_{t}^{\theta}$ to (4.1).
Proof.
The boundedness of $\theta_{t}$ ensures that the modified diffusion coefficient $\sigma(t,x)\theta_{t}$ satisfies:
$|\sigma(t,x)\theta_{t}-\sigma(t,y)\theta_{t}|$ $\leq M|\sigma(t,x)-\sigma(t,y)|\leq MK|x-y|$
$|\sigma(t,x)\theta_{t}|$ $\leq M|\sigma(t,x)|\leq MK(1+|x|)$
Thus the Lipschitz condition (4.3) and growth condition (4.4) hold for the controlled SDE. Existence and uniqueness follow from standard SDE theory. ∎
4.2 G-SDE Solution via Parameterization
Theorem 4.3 (G-SDE Solution).
Under the same Lipschitz and growth conditions, the G-SDE (4.2) admits a unique solution given by:
$X_{t}=\{X_{t}^{\theta}\}_{\theta\in\mathcal{A}^{\Theta}}$
where each $X_{t}^{\theta}$ solves the classical SDE (4.1).
Proof.
The solution is constructed in three steps:
  1. 1.
    For each $\theta\in\mathcal{A}^{\Theta}$, solve the classical SDE (4.1)
  2. 2.
    The boundedness of $\Theta$ ensures uniform estimates:
    $\sup_{\theta\in\mathcal{A}^{\Theta}}\mathbb{E}\left[\sup_{0\leq s\leq t}|X_{s}^{\theta}|^{2}\right]\leq C(1+|X_{0}|^{2})e^{Ct}$
  3. 3.
    The solution family $\{X_{t}^{\theta}\}$ satisfies the consistency conditions for G-expectation
The compactness of $\Theta$ is crucial for preserving the solution properties:
[…]
The boundedness of the control set $\Theta$ is essential for:
  • Preserving the Lipschitz and growth conditions under parameterization
  • Ensuring existence and uniqueness of solutions
  • Maintaining moment estimates and path regularity
This framework provides a rigorous bridge from classical SDE theory to G-SDEs, leveraging the bounded control structure to transfer solution properties.
The G-SDE solution is reconstructed from the parameterized solutions:
[…]
The solution inherits properties from the parameterized solutions:
Theorem 4.4 (Properties of Solutions ).
The G-SDE solution satisfies:
  1. 1.
    Moment estimates: $\hat{\mathbb{E}}\left[\sup_{0\leq s\leq t}|X_{s}|^{p}\right]<\infty$ for $p\geq 2$
  2. 2.
    Continuous dependence: If $b_{n}\to b$, $\sigma_{n}\to\sigma$ uniformly, then $X_{t}^{n}\to X_{t}$ in $L_{G}^{2}$
4.3 Geometric Brownian Motion — A Clear Example
Consider the geometric-type G-SDE with volatility uncertainty:
$dX_{t}=\sigma X_{t}dB_{t}+\gamma X_{t}d\langle B\rangle_{t},\quad X_{0}=x_{0}>0$ (4.6)
Through the isomorphism $\Phi$, we transform the G-SDE into a family of classical SDEs:
$\Phi(X_{t})=\{X_{t}^{\theta}\}_{\theta\in\mathcal{A}^{\Theta}}$
where for each control process $\theta\in\mathcal{A}^{\Theta}$
$dX_{t}^{\theta}=\sigma X_{t}^{\theta}\theta_{t}dW_{t}+\gamma X_{t}^{\theta}\theta_{t}^{2}dt,\quad X_{0}^{\theta}=x_{0}$ (4.7)
Applying the G-Itô formula to $f(x)=\ln x$:
$d(\ln X_{t})$ $=\sigma dB_{t}+\gamma d\langle B\rangle_{t}-\frac{1}{2}\sigma^{2}d\langle B\rangle_{t}$ (4.8)
$=\sigma dB_{t}+\left(\gamma-\frac{1}{2}\sigma^{2}\right)d\langle B\rangle_{t}$ (4.9)
Integrating yields the solution:
$X_{t}=x_{0}\exp\left(\sigma B_{t}+\left(\gamma-\frac{1}{2}\sigma^{2}\right)\langle B\rangle_{t}\right)$ (4.10)
For each fixed control $\theta$, apply classical Itô’s formula:
$d(\ln X_{t}^{\theta})$ $=\sigma\theta_{t}dW_{t}+\gamma\theta_{t}^{2}dt-\frac{1}{2}\sigma^{2}\theta_{t}^{2}dt$ (4.11)
$=\sigma\theta_{t}dW_{t}+\left(\gamma-\frac{1}{2}\sigma^{2}\right)\theta_{t}^{2}dt$ (4.12)
Integrating gives the parameterized solution:
$X_{t}^{\theta}=x_{0}\exp\left(\sigma\int_{0}^{t}\theta_{s}dW_{s}+\left(\gamma-\frac{1}{2}\sigma^{2}\right)\int_{0}^{t}\theta_{s}^{2}ds\right)$ (4.13)
Note that the two are identical under isomorphism.
The G-expectation is obtained by taking supremum over controls:
$\hat{\mathbb{E}}[X_{t}]=\sup_{\theta\in\mathcal{A}^{\Theta}}\mathbb{E}_{\mathbb{P}}[X_{t}^{\theta}]$ (4.14)
Thus the G-expectation is:
$\hat{\mathbb{E}}[X_{t}]=\sup_{\theta\in[\underline{\sigma},\overline{\sigma}]}x_{0}\exp(\gamma\theta^{2}t)=\begin{cases}x_{0}\exp(\gamma\overline{\sigma}^{2}t),&\gamma>0\\ x_{0}\exp(\gamma\underline{\sigma}^{2}t),&\gamma<0\end{cases}$ (4.15)
5 Properties of G-Martingales
In G-expectation theory, martingales exhibit fundamentally different behaviors due to volatility uncertainty. This chapter explains symmetric and asymmetric martingales through the lens of parameterized controls, revealing their economic interpretations as different types of ”fair gambles”.
Under the isomorphism $\Phi$, a G-martingale $M_{t}$ is represented as a family of classical processes:
$\Phi(M_{t})=\{M_{t}^{\theta}\}_{\theta\in\mathcal{A}^{\Theta}}$
where each $M_{t}^{\theta}$ is adapted to the filtration generated by $W_{t}^{\theta}$.
5.1 Symmetric Martingales: Fair Gambles in All Scenarios
Definition 5.1 (Symmetric Martingale).
A process $M_{t}$ is a symmetric G-martingale if for every control $\theta\in\mathcal{A}_{s,t}^{\Theta}$, $M_{t}^{\theta}$ is a classical martingale under $\mathbb{P}$:
$\mathbb{E}_{\mathbb{P}}[M_{t}^{\theta}|\mathcal{F}_{s}]=M_{s}^{\theta}\quad\forall s\leq t.$
[…]
5.2 Asymmetric Martingales: Favorable Gambles with Worst-Case Fairness
Definition 5.2 (Asymmetric Martingale).
A process $M_{t}$ is an asymmetric G-martingale if:
  1. 1.
    For all $\theta\in\mathcal{A}_{s,t}^{\Theta}$, $M_{t}^{\theta}$ is a supermartingale under $\mathbb{P}$:
    $\mathbb{E}_{\mathbb{P}}[M_{t}^{\theta}|\mathcal{F}_{s}]\leq M_{s}^{\theta}\quad\forall s\leq t$
  2. 2.
    There exists a worst-case scenario $\theta^{*}\in\Theta$ such that:
    $\mathbb{E}_{\mathbb{P}}[M_{t}^{\theta^{*}}|\mathcal{F}_{s}]=M_{s}^{\theta^{*}}$
5.3 Key Properties
[…]
6 Conclusion
This paper has established a fundamental and rigorous connection between the abstract theory of $G$-expectation and the concrete framework of parameterized stochastic control. By introducing the concept of controlled processes $\{B_{t}^{\theta}\}_{\theta\in\mathcal{A}^{\Theta}}$, we have constructed a complete isomorphism $\Phi$ that maps the nonlinear $G$-expectation space onto a family of classical probability spaces indexed by volatility controls.
The core contribution of this work is threefold:
  1. 1.
    Unified Representation: We demonstrated that $G$-Brownian motion and $G$-SDE solutions admit a natural representation as suprema over parameterized classical processes, providing an intuitive and computationally tractable interpretation of model uncertainty.
  2. 2.
    Algebraic and Analytic Isomorphism: The mapping $\Phi$ was shown to preserve all essential operations – addition, multiplication, function composition, and limit transitions – thereby maintaining the complete mathematical structure while transitioning between nonlinear and classical frameworks.
  3. 3.
    Martingale Classification: Our framework provided a clear distinction between symmetric $G$-martingales (fair in all volatility scenarios) and asymmetric $G$-martingales (favorable but fair only in worst-case scenarios), revealing the fundamental structure of stochastic processes under volatility uncertainty.
The parameterized control perspective developed herein offers more than theoretical elegance; it provides a practical computational methodology for working with $G$-expectations by reducing complex nonlinear problems to families of more manageable classical problems. This approach maintains full mathematical rigor while offering significant advantages for both theoretical analysis and numerical implementation.
Looking forward, this isomorphism establishes a solid foundation for further research in nonlinear stochastic calculus, with particular promise for applications in mathematical finance where volatility uncertainty plays a crucial role in pricing and risk management. The framework naturally extends to more complex stochastic processes and offers a new paradigm for understanding model uncertainty in quantitative economics.